OpenStudy (mayankdevnani):

Solve this ONE :- Q-) Problem on SERIES. 6,20,36,48,50,?,0 a)36 b)40 c)46 d)56 Which number comes in place of "?" ?

OpenStudy (mayankdevnani):

@ash2326 @phi @UsukiDoll

OpenStudy (anonymous):

56?

OpenStudy (mayankdevnani):

how?

OpenStudy (usukidoll):

Ned Flanders

OpenStudy (mayankdevnani):

what?

OpenStudy (mayankdevnani):

@INeedHelpPlease? help me

OpenStudy (raden):

Tn = 7n^2 - n^3 T1 = 7(1)^2 - (1)^3 = 6 T2 = 7(2)^2 - (2)^3 = 20 .... ... T6 = 7(6)^2 - (6)^3 = 78 T7 = 7(7)^2 - (7)^3 = 0 so, it is T6 = 78

OpenStudy (mayankdevnani):

but it is not in options. @RadEn

OpenStudy (mayankdevnani):

i know this is also a answer but i need answer which is in options?

OpenStudy (mayankdevnani):

@hartnn can you help me ?

hartnn (hartnn):

how did u get Tn RadEn ?

OpenStudy (anonymous):

@RadEn's reasoning checks out, his answer should be 36 though. \(7(6^2)-6^3=252-216=36\)

OpenStudy (mayankdevnani):

he makes the series to solve it. @hartnn

hartnn (hartnn):

yeah, i know how to find Tn, but i wanted to know whether RadEn used any different/simpler method....

OpenStudy (mayankdevnani):

his method is wrong.

hartnn (hartnn):

he got the answer which is in the choices, how u say its wrong ?

OpenStudy (mayankdevnani):

because if we put n=6,the answer comes to be is 36 but in actual(followed by series)the answer would be 50

OpenStudy (mayankdevnani):

isiliye mujhe lagta h ki voh galat h

hartnn (hartnn):

" in actual(followed by series)the answer would be 50" what does this even mean ?

OpenStudy (mayankdevnani):

iska matla agar hum uske tarike se follow karien to iska answer 50 aana chaihye series ke hisaab se lekin actual main 36 aa raha h..... iska method thode se hi digits ke liye apply hota h

OpenStudy (mayankdevnani):

now understood @hartnn bhaiya

hartnn (hartnn):

lol, i could translate that in hindi myself too

hartnn (hartnn):

why u say answer should be 50 ?

OpenStudy (mayankdevnani):

bhaiya , yeh question ntse 2013 RAJ. main poocha gaya tha aur FIITJEE vaalein bata rahein h ki iska answer DATA INADEQUATE h aur CAREER POINT vallein bata rahein h ki iska answer 36 h

OpenStudy (mayankdevnani):

to iska kya answer hoga bhaiya?

OpenStudy (mayankdevnani):

@hartnn

hartnn (hartnn):

the answer to this is indeed 36 as shown corretly by RadEn

OpenStudy (mayankdevnani):

okk...thank you,,, @hartnn and @RadEn

hartnn (hartnn):

i just wanted to know any simpler way of finding Tn :P

OpenStudy (mayankdevnani):

yeah...this logic doesn't apply in NTSE STAGE 1

OpenStudy (mayankdevnani):

i think there is another logic simpler logic for this

hartnn (hartnn):

i think it does apply, the third differences are constant

hartnn (hartnn):